I need help to solve X in
square root x+7 equals x+1
thanks

Answer by venugopalramana(3286)

square root x+7 equals x+1 ..SQUARE BOTH SIDES
X+7=(X+1)^2=X^2+2X+1
X^2+2X+1-X-7=0
X^2+X-6=0
X^2+3X-2X-6=0
X(X+3)-2(X+3)=0
(X-2)(X+3)=0
X-2=0...THAT IS...X=2....OR......X+3=0...THAT IS X=-3